OpenStudy (anonymous):

Use the Fundamental Theorem of Algebra to find the number of complex roots of the polynomial, including any repeated roots. x3+3x+17

OpenStudy (anonymous):

@wolfe8

OpenStudy (anonymous):

Theres no multiple choice?? lmao

OpenStudy (anonymous):

I don't have multiple choices

OpenStudy (anonymous):

Oh ill help you on the next one i know the formula i just dont know how to do the problem lol

OpenStudy (wolfe8):

Ok I'm not home and I used a software for this so this is what I have for now until I get home and do it myself: Change coordinates by substituting x = lambda/y+y, where lambda is a constant value that will be determined later. If x = lambda/y+y then y = 1/2 (sqrt(x^2-4 lambda)+x) which will be used during back substitution: 17+3 (y+lambda/y)+(y+lambda/y)^3 = 0 Transform the rational equation into a polynomial equation. Multiply both sides by y^3 and collect in terms of y: 17 y^3+y^6+lambda^3+y^4 (3+3 lambda)+y^2 (3 lambda+3 lambda^2) = 0 Find an appropriate value for lambda in order to make the coefficients of y^2 and y^4 both zero. Substitute lambda = -1 and then z = y^3, yielding a quadratic equation in the variable z: -1+17 z+z^2 = 0 Solve for z. Find the positive solution to the quadratic equation: z = 1/2 (-17+sqrt(293)) Perform back substitution on z = 1/2 (sqrt(293)-17). Substitute back for z = y^3: y^3 = 1/2 (-17+sqrt(293)) Take the cube root of both sides. Taking cube roots gives (1/2 (sqrt(293)-17))^(1/3) times the third roots of unity: y = (1/2 (-17+sqrt(293)))^(1/3) or y = -(1/2 (17-sqrt(293)))^(1/3) or y = (-1)^(2/3) (1/2 (-17+sqrt(293)))^(1/3) Look at the first equation: Perform back substitution on y = (1/2 (sqrt(293)-17))^(1/3). Substitute back for y = sqrt(x^2+4)/2+x/2: x/2+sqrt(4+x^2)/2 = (1/2 (-17+sqrt(293)))^(1/3) or y = -(1/2 (17-sqrt(293)))^(1/3) or y = (-1)^(2/3) (1/2 (-17+sqrt(293)))^(1/3) Solve for x. Solve the radical equation in x: x = -(2/(-17+sqrt(293)))^(1/3)+(1/2 (-17+sqrt(293)))^(1/3) or y = -(1/2 (17-sqrt(293)))^(1/3) or y = (-1)^(2/3) (1/2 (-17+sqrt(293)))^(1/3) Look at the second equation: Perform back substitution on y = -(1/2 (17-sqrt(293)))^(1/3). Substitute back for y = sqrt(x^2+4)/2+x/2: x = -(2/(-17+sqrt(293)))^(1/3)+(1/2 (-17+sqrt(293)))^(1/3) or x/2+sqrt(4+x^2)/2 = -(1/2 (17-sqrt(293)))^(1/3) or y = (-1)^(2/3) (1/2 (-17+sqrt(293)))^(1/3) Solve for x. Solve the radical equation in x: x = -(2/(-17+sqrt(293)))^(1/3)+(1/2 (-17+sqrt(293)))^(1/3) or x = -(1/2 (17-sqrt(293)))^(1/3)-(-1)^(2/3) (2/(-17+sqrt(293)))^(1/3) or y = (-1)^(2/3) (1/2 (-17+sqrt(293)))^(1/3) Look at the third equation: Perform back substitution on y = (-1)^(2/3) (1/2 (sqrt(293)-17))^(1/3). Substitute back for y = sqrt(x^2+4)/2+x/2: x = -(2/(-17+sqrt(293)))^(1/3)+(1/2 (-17+sqrt(293)))^(1/3) or x = -(1/2 (17-sqrt(293)))^(1/3)-(-1)^(2/3) (2/(-17+sqrt(293)))^(1/3) or x/2+sqrt(4+x^2)/2 = (-1)^(2/3) (1/2 (-17+sqrt(293)))^(1/3) Solve for x. Solve the radical equation in x: Answer: | | x = -(2/(-17+sqrt(293)))^(1/3)+(1/2 (-17+sqrt(293)))^(1/3) or x = -(1/2 (17-sqrt(293)))^(1/3)-(-1)^(2/3) (2/(-17+sqrt(293)))^(1/3) or x = (-2/(-17+sqrt(293)))^(1/3)+(-1)^(2/3) (1/2 (-17+sqrt(293)))^(1/3)

OpenStudy (anonymous):

Great! Thank you so much!

OpenStudy (petiteme):

This is insane lol

OpenStudy (anonymous):

Hahaha

OpenStudy (whpalmer4):

The problem asks you to find the NUMBER of complex roots, does it not?

OpenStudy (whpalmer4):

The equation \[x^3+3x+17=0\] has how many roots in total?

OpenStudy (whpalmer4):

One of the corollaries of the FToA is that every polynomial in a single variable x with real coefficients (no \(i\) or \(\sqrt{-1}\) anywhere) can be written as a product of a constant, factors such as \((x+a)\) where \(a\) is real, and factors such as \((x^2+ax + b)\) where \(a,b\) are real and \(a^2-4b <0\), which means \((x^2+ax+b)\) has only complex roots.

OpenStudy (whpalmer4):

If you can figure out how many roots the equation has (which you can, just by looking at it and observing the highest power of \(x\)) and recalling that a polynomial in one variable with only real coefficients must have an even number of complex roots (because they come in conjugate pairs), can you conclude anything about this problem?